Find the degree measure of an angle which is 4/5 of its supplement

Answer by cleomenius(959)

You will have 9 parts total.
180 / 9 = 20
4 * 20 = 80 degrees
5 * 20 = 100 degrees.
80/100 = 4/5.
Cleomenius.